MF Corp. has an ROE of 17% and a plowback ratio of 55%. The market capitalization rate is 15%.

a. If the coming year’s earnings are expected to be $2.10 per share, at what price will the stock sell? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. What price do you expect MF shares to sell for in five years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1
Growth rate=ROE*retention ratio
growth rate=17*0.55
growth rate = 9.35

a

Dividend = EPS*(1-plowback) = 2.1*(1-0.55) = 0.945

Price = Dividend in 1 year/(cost of equity - growth rate)
Price = 0.945/ (0.15 - 0.0935)
Price = 16.73

b

Price in 5 year = current price*(1+cap rate-growth rate)^5

=16.73*(1+0.15-0.0935)^5 = 22.02